# Agent-X: Evaluating Deep Multimodal Reasoning in Vision-Centric Agentic Tasks

### v1

@@ F.3 Reasoning Chain Experiments
Appendix G Instruction for Annotators
@@
Appendix H Agent-X Task Examples
@@
Example 5 Refer to caption AgentX_133.jpg Query: What is the solution to the equation, and is the solution real, complex, or integer? Involved Tools: MathOCR, Solver, Calculator Steps: 1. Extract the mathematical expression from the image. 2. Solve the extracted equation to find the values of x 3. Analyze the solutions to determine whether they are real, complex, or integers. Reasoning Trace for steps: 1. The MathOCR tool extracts the equation $(x-sqrt(2))^{2}+4*sqrt(2)*x=0$ from the image, enabling symbolic processing. 2. The Solver tool is used to solve the equation, yielding the solution $x=-sqrt(2)$. 3. The Calculator confirms that the value $x=-sqrt(2)$ is a real number, which is irrational but not complex or integer. Answer: $x=-sqrt(2)$. Real Justification: The equation was extracted, solved, and evaluated. There is just one solution of this equation. The resulting value is real and irrational, not complex or integer.
@@
Appendix I Generation Prompts
@@
Figure 23: Prompt given to GPT4-o to for the initial query generation
Reasoning and Final Answer "You are an agent performing multimodal reasoning using the tools listed below.
Your goal is to answer the given query step-by-step by selecting the right tools for each stage of reasoning.
For each step: Clearly state what you’re trying to do. Specify the tool you are using. Explicitly include the input provided to the tool and the output received from the tool. Explain what you learned or why this step was necessary. Make sure your tool usage follows the tool descriptions provided. Do not hallucinate tools or skip intermediate reasoning steps. --- Available Tools & How to Use Them: {toolmeta_section} --- Query: "{query}" Images: {image1, image2, ...} --- Return your output as a JSON list of steps, formatted like this: [ {
"step": <step number>,
"task": "<Short description of what this step is trying to do>",
"tool": "<Tool name(s)>",
"input": "<Input provided to the tool>",
"output": "<Output from the tool>",
"thought": "<Why this step was needed, or what you learned>"
"final_answer": {
"value": "<Final conclusion>",
"justification": "<How all steps lead to the answer>"
Figure 24: Prompt given to GPT4-o for the initial reasoning and final answer generation
Appendix J Evaluation Prompts
Grounding Score You are an evaluation assistant measuring the groundedness of each reasoning step performed by a vision-language agent.
"You are given a Ground Truth (GT) block containing the original query, a sequence of reasoning steps, and the final answer along with its justification. Each reasoning step includes the task being attempted, the tool used (with its input and output), and the agent’s thought process.
You are also given the full reasoning trace produced by the agent, including each step’s task, selected tool and its output, and thought.
Your task is to assess whether each agent step is visually and contextually grounded — that is, whether the task, the selected tool, the tool output, and the agent’s thought process are all supported by the GT.
Evaluation Guidelines: If the number of steps in the GT and the model differ: Any extra model step beyond the GT steps should receive a score of 0 (hallucinated or unjustified). Any GT step that the model omits should also receive a score of 0 (missing reasoning). The Grounding Score for the full query is computed as the average of the per-step scores. Assess the following per step: Is the task relevant to the query and aligned with the GT? Is the tool appropriate for this task? Is the output consistent with what’s in the GT? Is the thought grounded and logically aligned? Scoring Criteria: 1 — All aspects are grounded (task, tool, output, and thought) 0.5 — Partially grounded (some align, some don’t) 0 — Ungrounded or hallucinated Output format:
Score: <0, 0.5, or 1>
Justification: <1--2 sentence explanation for the score>"
@@
Tool Precision "You are an evaluation assistant measuring tool precision in an agent’s reasoning step.
You are given a Ground Truth (GT) block containing the original query, a sequence of reasoning steps, and the final answer along with its justification. Each reasoning step includes the task being attempted, the tool used (with its input and output), and the agent’s thought process.
You are also given the full reasoning trace produced by the agent, including each step’s task, selected tool and its output, and thought.
Your task is to assess whether the tool selected by the agent in each reasoning step is the most appropriate tool for the task, by comparing it with the GT.
Evaluation Guidelines: If the number of steps in the GT and the model differ: Any extra model step beyond the GT steps should receive a score of 0 (hallucinated or unjustified). Any GT step that the model omits should also receive a score of 0 (missing reasoning). The Tool Precision score for the full query is computed as the average of the per-step scores. Assess the following per step: Is the tool selected by the agent the same as the one used in the GT? Is the tool the most appropriate for the task being attempted? Scoring Criteria: 1 — The selected tool matches the GT tool and is appropriate for the task. 0 — The tool does not match the GT or is an inappropriate choice. Output format:
Score: <0 or 1>
Justification: <1--2 sentence explanation for the score>"
@@
Tool Accuracy "You are an evaluation assistant measuring tool accuracy in an agent’s reasoning step.
You are given the task the agent is trying to accomplish, the tool it uses along with its input and output, and the tool metadata containing a description of the tool’s purpose and its expected input/output formats.
Your task is to assess whether the tool used by the agent was applied correctly in each step by checking: Whether the output format is valid and consistent with the tool’s specification (based on tool metadata). Whether the output is relevant and meaningful for completing the step’s task. Evaluation Guidelines: The Tool Accuracy score for the full query is computed as the average of the per-step scores. Assess the following per step: Is the tool’s output correctly formatted according to the tool metadata? Is the output meaningful and appropriate for the stated task? Scoring Criteria: 1 — Output is valid, properly formatted, and clearly relevant to the task. 0.5 — Output is partially valid or partially relevant. 0 — Output is incorrectly formatted, irrelevant, or unhelpful. Output format:
Score: <0, 0.5, or 1>
Justification: <1--2 sentence explanation for the score>"
@@
Faithfulness Accuracy "You are an evaluation assistant measuring the Faithfulness Accuracy of an agent’s reasoning process.
You are given a Ground Truth (GT) block containing the original query, a sequence of reasoning steps, and the final answer along with its justification. Each reasoning step includes the task being attempted, the tool used (with its input and output), and the agent’s thought process.
You are also given the full reasoning trace produced by the agent, including each step’s task, selected tool, output, and thought.
Your task is to assess how faithful the agent’s reasoning trace is to the GT, i.e., whether the steps follow a logically sound plan that aligns with the structure, intent, and direction of the GT. Evaluation Guidelines: Focus on the structure and logical flow of the agent’s reasoning steps. Determine whether the steps collectively form a coherent strategy to answer the query. Faithfulness is about consistency with the GT’s method, not necessarily correctness of individual steps. Scoring Criteria: 1 — The reasoning is faithful to the GT: it follows a logically sound plan that mirrors the GT in structure and direction. 0.5 — The reasoning partially follows the GT’s structure, but contains some deviations, redundancies, or inconsistencies. 0 — The reasoning is not faithful to the GT and lacks logical progression or alignment with the intended plan. Output format:
Score: <0, 0.5, or 1>
Justification: <1--2 sentence explanation for the score>"
@@
Context Score "You are an evaluation assistant measuring the Context Score of an agent’s reasoning step.
You are given a Ground Truth (GT) block containing the original query, a sequence of reasoning steps, and the final answer along with its justification. Each reasoning step includes the task being attempted, the tool used (with its input and output), and the agent’s thought process.
You are also given the full reasoning trace produced by the agent, including each step’s task, selected tool and its output, and thought, as well as the agent’s final answer and justification.
Your task is to assess whether the agent’s reasoning step is grounded in the input context, and whether that context was effectively used in the reasoning process. Inputs may include image, video, text, audio, or a combination. Evaluation Guidelines: If the number of steps in the GT and the model differ: Any extra model step beyond the GT steps should receive a score of 0 (hallucinated or unjustified). Any GT step that the model omits should also receive a score of 0 (missing reasoning). The Context Score for the full query is computed as the average of the per-step scores. Assess the following per step: Does the agent correctly use relevant parts of the input? Is the input used appropriate for the tool selected and the task being attempted? Does the reasoning show effective and meaningful use of the input? Scoring Criteria: 1 — The agent uses the input fully and appropriately. 0.5 — The agent uses some relevant input but misses important details or misuses others. 0 — The agent ignores or misinterprets the input entirely. Output format:
Score: <0, 0.5, or 1>
Justification: <1--2 sentence explanation for the score>"
@@
Factual Accuracy "You are an evaluation assistant measuring the factual accuracy of the agent’s reasoning steps.
You are given a Ground Truth (GT) block containing the original query, a sequence of reasoning steps, and the final answer along with its justification. Each reasoning step includes the task being attempted, the tool used (with its input and output), and the agent’s thought process.
You are also given the full reasoning trace produced by the agent, including each step’s task, selected tool, tool input/output, and thought.
Your task is to assess whether the agent introduces any hallucinated, fabricated, or factually incorrect information in its reasoning when compared to the GT. Evaluation Guidelines: Compare the model’s reasoning steps to the GT to identify factual hallucinations or incorrect claims. Focus on whether the output or thought includes details not supported by the GT. Do not penalize for minor omissions unless they lead to a factual error. Scoring Criteria: 1 — No factual errors or hallucinations compared to the GT. 0.5 — Minor inaccuracies or vague mismatches. 0 — Major factual errors or hallucinated content. Output format:
Score: <0, 0.5, or 1>
Justification: <1--2 sentence explanation for the score>"
@@
Semantic Accuracy "You are an evaluation assistant measuring the Semantic Accuracy of an agent’s reasoning process.
You are given a Ground Truth (GT) block containing the original query, a sequence of reasoning steps, and the final answer along with its justification. Each reasoning step includes the task being attempted, the tool used (with its input and output), and the agent’s thought process.
You are also given the full reasoning trace produced by the agent, including each step’s task, selected tool and its output, and thought, as well as the agent’s final answer and justification.
@@
Your task is to assess whether the agent’s reasoning and final output semantically align with the Ground Truth, i.e., whether the agent has covered all the essential parts of the query as demonstrated in the GT. Evaluation Guidelines: Compare the agent’s reasoning trace and final answer with the GT to check whether all key components of the query are addressed. Credit should be given for meaningful semantic coverage, not superficial similarity. If the model ignores or misunderstands core parts of the GT reasoning or final answer, penalize accordingly. Scoring Criteria: 1 — The agent addresses all key components of the query, matching the GT’s semantic scope. 0.5 — Some parts are covered, but the response misses or weakly handles other essential elements. 0 — The agent’s reasoning or answer omits or misrepresents major parts of the query. Output format:
Score: <0, 0.5, or 1>
Justification: <1--2 sentence explanation for the score>"
@@
Figure 31: Prompt for evaluating semantic consistency of the agent’s reasoning and answer
Goal Accuracy "You are an evaluation assistant measuring the Goal Accuracy of a vision-language agent’s final answer.
You are given the original query, the agent’s final output, the ground truth (GT) final answer, and the type of query — either ‘‘objective’’ or ‘‘subjective’’.
@@
Your task is to assess how well the agent’s final output matches the ground truth answer, based on the nature of the query. Evaluation Guidelines: The Goal Accuracy score is computed once per query (not per step). Use exact match evaluation for objective queries. Use semantic similarity evaluation for subjective queries. Scoring Criteria: For objective queries: 1 — The final output matches the GT exactly or is clearly equivalent. 0 — The output is incorrect, incomplete, or unrelated. For subjective queries: Score = Cosine similarity between the agent’s answer and the GT answer (range: 0 to 1) Output format:
If objective:
Score: <0 or 1>
Justification: <Optional 1--2 sentence explanation>
If subjective:
Score: <cosine similarity>
Justification: <Optional 1--2 sentence explanation>"
@@
Figure 32: Prompt for evaluating alignment of final answer with the ground truth
Toolset Accuracy "You are an evaluation assistant measuring the toolset accuracy of an agent’s reasoning process.
You are given a Ground Truth (GT) block containing the original query, a sequence of reasoning steps, and the final answer along with its justification. Each reasoning step includes the task being attempted, the tool used (with its input and output), and the agent’s thought process.
You are also given the full reasoning trace produced by the agent, including each step’s task, selected tool and its output, and thought.
@@
Your task is to evaluate whether the agent used the correct tools overall by comparing the set of tools it used to the set used in the GT. Evaluation Guidelines: If the agent uses tools that are not present in the GT or misses tools that are, it should be penalized. The score reflects how well the agent’s toolset aligns with the GT toolset across the full reasoning trace. Output format:
Score: <F1 score rounded to 2 decimal places>
Justification: <1--2 sentence explanation for the score>"
@@
Figure 33: Prompt for evaluating alignment of the agent’s toolset with the ground truth
Appendix K Benchmark Curation Examples
@@
K.1 Reasoning Trace Generation
@@
K.2 Human Verification Tool

### v2

@@ G.6 Reasoning Chain Experiments
Figure 9: Tool Call analysis across LMM families on Agent-X.
Tool Call Success and Failure Analysis. Figure 9 shows the number of successful and failed tool invocations per model. A successful call executes without format or argument errors. The GPT family achieves the highest success rate (83.8%) relative to its total calls, with GPT-4o demonstrating a strong balance between usage and reliability. InternVL models lead to absolutely successful calls but exhibit high failure counts, primarily due to invalid tool usage. Qwen2.5-VL-8B shows the highest precision, with only 109 failures out of 2241 calls, reflecting robust format compliance. In contrast, o4-mini is the most aggressive, issuing 3374 calls but suffering the highest failure rate (1038). Gemini-3-4B and VideoLLaMA3-7B show moderate tool use, with Gemini-3-4B maintaining a lower failure count. These trends highlight three behavioral patterns: aggressive models ($e.g.$, InternVL3-8B) invoke tools frequently but are prone to errors; conservative models ($e.g.$, Qwen2.5-VL-7B) prioritize accuracy over volume; and balanced models ($e.g.$., GPT-4o) combine moderate tool use with high reliability.
Appendix H Instruction for Annotators
@@
Appendix I Agent-X Task Examples
@@
Example 5 Refer to caption AgentX_133.jpg Query: What is the solution to the equation, and is the solution real, complex, or integer? Involved Tools: MathOCR, Solver, Calculator Steps: 1. Extract the mathematical expression from the image. 2. Solve the extracted equation to find the values of x 3. Analyze the solutions to determine whether they are real, complex, or integers. Reasoning Trace for steps: 1. The MathOCR tool extracts the equation $(x-sqrt(2))^{2}+4*sqrt(2)*x=0$ from the image, enabling symbolic processing. 2. The Solver tool is used to solve the equation, yielding the solution x=sqrt(2)x=-sqrt(2). 3. The Calculator confirms that the value x=sqrt(2)x=-sqrt(2) is a real number, which is irrational but not complex or integer. Answer: x=sqrt(2)x=-sqrt(2). Real Justification: The equation was extracted, solved, and evaluated. There is just one solution of this equation. The resulting value is real and irrational, not complex or integer.
Figure 16: Example 5
@@
Appendix J Generation Prompts
@@
Figure 23: Prompt given to GPT4-o to for the initial query generation.
Reasoning and Final Answer "You are an agent performing multimodal reasoning using the tools listed below. Your goal is to answer the given query step-by-step by selecting the right tools for each stage of reasoning. For each step: Clearly state what you’re trying to do. Specify the tool you are using. Explicitly include the input provided to the tool and the output received from the tool. Explain what you learned or why this step was necessary. Make sure your tool usage follows the tool descriptions provided. Do not hallucinate tools or skip intermediate reasoning steps. --- Available Tools & How to Use Them: {toolmeta_section} --- Query: "{query}" Images: {image1, image2, ...} --- Return your output as a JSON list of steps, formatted like this: [ { "step": <step number>, "task": "<Short description of what this step is trying to do>", "tool": "<Tool name(s)>", "input": "<Input provided to the tool>", "output": "<Output from the tool>", "thought": "<Why this step was needed, or what you learned>" }, ... { "final_answer": { "value": "<Final conclusion>", "justification": "<How all steps lead to the answer>" } } ]"
Figure 24: Prompt given to GPT4-o for the initial reasoning and final answer generation.
Appendix K Evaluation Prompts
Grounding Score You are an evaluation assistant measuring the groundedness of each reasoning step performed by a vision-language agent. "You are given a Ground Truth (GT) block containing the original query, a sequence of reasoning steps, and the final answer along with its justification. Each reasoning step includes the task being attempted, the tool used (with its input and output), and the agent’s thought process. You are also given the full reasoning trace produced by the agent, including each step’s task, selected tool and its output, and thought. Your task is to assess whether each agent step is visually and contextually grounded — that is, whether the task, the selected tool, the tool output, and the agent’s thought process are all supported by the GT. Evaluation Guidelines: If the number of steps in the GT and the model differ: Any extra model step beyond the GT steps should receive a score of 0 (hallucinated or unjustified). Any GT step that the model omits should also receive a score of 0 (missing reasoning). The Grounding Score for the full query is computed as the average of the per-step scores. Assess the following per step: Is the task relevant to the query and aligned with the GT? Is the tool appropriate for this task? Is the output consistent with what’s in the GT? Is the thought grounded and logically aligned? Scoring Criteria: 1 — All aspects are grounded (task, tool, output, and thought) 0.5 — Partially grounded (some align, some don’t) 0 — Ungrounded or hallucinated Output format: Score: <0, 0.5, or 1> Justification: <1--2 sentence explanation for the score>"
@@
Tool Precision "You are an evaluation assistant measuring tool precision in an agent’s reasoning step. You are given a Ground Truth (GT) block containing the original query, a sequence of reasoning steps, and the final answer along with its justification. Each reasoning step includes the task being attempted, the tool used (with its input and output), and the agent’s thought process. You are also given the full reasoning trace produced by the agent, including each step’s task, selected tool and its output, and thought. Your task is to assess whether the tool selected by the agent in each reasoning step is the most appropriate tool for the task, by comparing it with the GT. Evaluation Guidelines: If the number of steps in the GT and the model differ: Any extra model step beyond the GT steps should receive a score of 0 (hallucinated or unjustified). Any GT step that the model omits should also receive a score of 0 (missing reasoning). The Tool Precision score for the full query is computed as the average of the per-step scores. Assess the following per step: Is the tool selected by the agent the same as the one used in the GT? Is the tool the most appropriate for the task being attempted? Scoring Criteria: 1 — The selected tool matches the GT tool and is appropriate for the task. 0 — The tool does not match the GT or is an inappropriate choice. Output format: Score: <0 or 1> Justification: <1--2 sentence explanation for the score>"
@@
Tool Accuracy "You are an evaluation assistant measuring tool accuracy in an agent’s reasoning step. You are given the task the agent is trying to accomplish, the tool it uses along with its input and output, and the tool metadata containing a description of the tool’s purpose and its expected input/output formats. Your task is to assess whether the tool used by the agent was applied correctly in each step by checking: Whether the output format is valid and consistent with the tool’s specification (based on tool metadata). Whether the output is relevant and meaningful for completing the step’s task. Evaluation Guidelines: The Tool Accuracy score for the full query is computed as the average of the per-step scores. Assess the following per step: Is the tool’s output correctly formatted according to the tool metadata? Is the output meaningful and appropriate for the stated task? Scoring Criteria: 1 — Output is valid, properly formatted, and clearly relevant to the task. 0.5 — Output is partially valid or partially relevant. 0 — Output is incorrectly formatted, irrelevant, or unhelpful. Output format: Score: <0, 0.5, or 1> Justification: <1--2 sentence explanation for the score>"
@@
Faithfulness Accuracy "You are an evaluation assistant measuring the Faithfulness Accuracy of an agent’s reasoning process. You are given a Ground Truth (GT) block containing the original query, a sequence of reasoning steps, and the final answer along with its justification. Each reasoning step includes the task being attempted, the tool used (with its input and output), and the agent’s thought process. You are also given the full reasoning trace produced by the agent, including each step’s task, selected tool, output, and thought. Your task is to assess how faithful the agent’s reasoning trace is to the GT, i.e., whether the steps follow a logically sound plan that aligns with the structure, intent, and direction of the GT. Evaluation Guidelines: Focus on the structure and logical flow of the agent’s reasoning steps. Determine whether the steps collectively form a coherent strategy to answer the query. Faithfulness is about consistency with the GT’s method, not necessarily correctness of individual steps. Scoring Criteria: 1 — The reasoning is faithful to the GT: it follows a logically sound plan that mirrors the GT in structure and direction. 0.5 — The reasoning partially follows the GT’s structure, but contains some deviations, redundancies, or inconsistencies. 0 — The reasoning is not faithful to the GT and lacks logical progression or alignment with the intended plan. Output format: Score: <0, 0.5, or 1> Justification: <1--2 sentence explanation for the score>"
@@
Context Score "You are an evaluation assistant measuring the Context Score of an agent’s reasoning step. You are given a Ground Truth (GT) block containing the original query, a sequence of reasoning steps, and the final answer along with its justification. Each reasoning step includes the task being attempted, the tool used (with its input and output), and the agent’s thought process. You are also given the full reasoning trace produced by the agent, including each step’s task, selected tool and its output, and thought, as well as the agent’s final answer and justification. Your task is to assess whether the agent’s reasoning step is grounded in the input context, and whether that context was effectively used in the reasoning process. Inputs may include image, video, text, audio, or a combination. Evaluation Guidelines: If the number of steps in the GT and the model differ: Any extra model step beyond the GT steps should receive a score of 0 (hallucinated or unjustified). Any GT step that the model omits should also receive a score of 0 (missing reasoning). The Context Score for the full query is computed as the average of the per-step scores. Assess the following per step: Does the agent correctly use relevant parts of the input? Is the input used appropriate for the tool selected and the task being attempted? Does the reasoning show effective and meaningful use of the input? Scoring Criteria: 1 — The agent uses the input fully and appropriately. 0.5 — The agent uses some relevant input but misses important details or misuses others. 0 — The agent ignores or misinterprets the input entirely. Output format: Score: <0, 0.5, or 1> Justification: <1--2 sentence explanation for the score>"
@@
Factual Accuracy "You are an evaluation assistant measuring the factual accuracy of the agent’s reasoning steps. You are given a Ground Truth (GT) block containing the original query, a sequence of reasoning steps, and the final answer along with its justification. Each reasoning step includes the task being attempted, the tool used (with its input and output), and the agent’s thought process. You are also given the full reasoning trace produced by the agent, including each step’s task, selected tool, tool input/output, and thought. Your task is to assess whether the agent introduces any hallucinated, fabricated, or factually incorrect information in its reasoning when compared to the GT. Evaluation Guidelines: Compare the model’s reasoning steps to the GT to identify factual hallucinations or incorrect claims. Focus on whether the output or thought includes details not supported by the GT. Do not penalize for minor omissions unless they lead to a factual error. Scoring Criteria: 1 — No factual errors or hallucinations compared to the GT. 0.5 — Minor inaccuracies or vague mismatches. 0 — Major factual errors or hallucinated content. Output format: Score: <0, 0.5, or 1> Justification: <1--2 sentence explanation for the score>"
@@
Semantic Accuracy "You are an evaluation assistant measuring the Semantic Accuracy of an agent’s reasoning process. You are given a Ground Truth (GT) block containing the original query, a sequence of reasoning steps, and the final answer along with its justification. Each reasoning step includes the task being attempted, the tool used (with its input and output), and the agent’s thought process. You are also given the full reasoning trace produced by the agent, including each step’s task, selected tool and its output, and thought, as well as the agent’s final answer and justification. Your task is to assess whether the agent’s reasoning and final output semantically align with the Ground Truth, i.e., whether the agent has covered all the essential parts of the query as demonstrated in the GT. Evaluation Guidelines: Compare the agent’s reasoning trace and final answer with the GT to check whether all key components of the query are addressed. Credit should be given for meaningful semantic coverage, not superficial similarity. If the model ignores or misunderstands core parts of the GT reasoning or final answer, penalize accordingly. Scoring Criteria: 1 — The agent addresses all key components of the query, matching the GT’s semantic scope. 0.5 — Some parts are covered, but the response misses or weakly handles other essential elements. 0 — The agent’s reasoning or answer omits or misrepresents major parts of the query. Output format: Score: <0, 0.5, or 1> Justification: <1--2 sentence explanation for the score>"
@@
Figure 31: Prompt for evaluating semantic consistency of the agent’s reasoning and answer.
@@
Goal Accuracy "You are an evaluation assistant measuring the Goal Accuracy of a vision-language agent’s final answer. You are given the original query, the agent’s final output, the ground truth (GT) final answer, and the type of query — either ‘‘objective’’ or ‘‘subjective’’. Your task is to assess how well the agent’s final output matches the ground truth answer, based on the nature of the query. Evaluation Guidelines: The Goal Accuracy score is computed once per query (not per step). Use exact match evaluation for objective queries. Use semantic similarity evaluation for subjective queries. Scoring Criteria: For objective queries: 1 — The final output matches the GT exactly or is clearly equivalent. 0 — The output is incorrect, incomplete, or unrelated. For subjective queries: Score = Cosine similarity between the agent’s answer and the GT answer (range: 0 to 1) Output format: If objective: Score: <0 or 1> Justification: <Optional 1--2 sentence explanation> If subjective: Score: <cosine similarity> Justification: <Optional 1--2 sentence explanation>"
@@
Figure 32: Prompt for evaluating alignment of final answer with the ground truth.
@@
Toolset Accuracy "You are an evaluation assistant measuring the toolset accuracy of an agent’s reasoning process. You are given a Ground Truth (GT) block containing the original query, a sequence of reasoning steps, and the final answer along with its justification. Each reasoning step includes the task being attempted, the tool used (with its input and output), and the agent’s thought process. You are also given the full reasoning trace produced by the agent, including each step’s task, selected tool and its output, and thought. Your task is to evaluate whether the agent used the correct tools overall by comparing the set of tools it used to the set used in the GT. Evaluation Guidelines: If the agent uses tools that are not present in the GT or misses tools that are, it should be penalized. The score reflects how well the agent’s toolset aligns with the GT toolset across the full reasoning trace. Output format: Score: <F1 score rounded to 2 decimal places> Justification: <1--2 sentence explanation for the score>"
@@
Figure 33: Prompt for evaluating alignment of the agent’s toolset with the ground truth.
Appendix L Benchmark Curation Examples
@@
L.1 Reasoning Trace Generation
@@
L.2 Human Verification Tool